• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name sawatzky

    Log entry time 00:32:46 on June 07, 2005

    Entry number 145383

    keyword=Shift summary: Monday swing

    Shift Crew:  Sawatzky, Perdrisat, Voutier


    Events:
    16:00 Currently in controlled access for polarity change. Eugene working on RHRS.

    17:20 Polarity on Left dipole and septum changed to positive

    Arun took a look at the BCM equipment and found the temperature control for the unser box had tripped. The temp takes ~24 hours to stabilize so the BCM calibration scheduled for Wed morning (may) get pushed back.

    18:00 Sieves out on both arms

    18:40 Hall A going to beam permit

    22:10 Test beam to Hall A

    22:50 Septum steering test begun

    23:55 Septum steering test completed: slight movement observed on beam dump viewer, but no steering adjustments were made to keep it 'on screen'


    Run Summary:
    No production runs.